5. Using the following data, perform a oneway analysis of variance using α=0.05.

group 1 group 2 group 3

51 23 56

45 43 76

33 23 74

45 43 87

67 45 56

a) What are the null and alternative hypotheses for this test?

b) What is your conclusion?
Answer #1

Overall total = 767
Overall mean Y̅.. = 767 / 15 = 51.1333

Y̅ .. is overall mean
y̅i . is treatment mean
SS total = ΣΣ(Yij - & Y̅..)2 = 4883.7332
SS treatment = ΣΣ(Yij - & y̅i.)2 = 3022.9332
SS error = Σ(y̅i. - & Y̅..)2 = 1860.8

MS treatment = ΣΣ(Yij - & y̅i.)2 / a - 1 = 1511.4666
MS error = Σ(y̅i. - & Y̅..)2 / N - a = 155.0667

ANOVA table
Source SS    df MS F    p-value
Treatment 3,022.93 2 1,511.467 9.75 0.0031
Error 1,860.80 12 155.067
Total 4,883.73 14

To Test :-
H0 :- µ1 = µ2 = µ3 = 0 i.e There is no difference in the group means
H0 :- µ1 = µ2 = µ3 ≠ 0 i.e Some means are different

Test Statistic :-
f = MS treatment / MS error = 9.7472

Test Criteria :-
Reject null hypothesis if f > f(α , a-1 , N-a )
Critical value f(0.05, 2 , 12 ) = 3.8853 (From F table)
Since 9.7472 > 3.8853, we reject H0
conclusion = Group means may differs